OpenStudy (anonymous):

f(x) = x3 – 2x2 + 16x – 32

mathslover (mathslover):

Hello mellali94, please specify your question. What do you want to do with the equation?

OpenStudy (solomonzelman):

If you need roots/zeros, factor the y-intercept, is always the only constant. @mellali94 do what mathsolver asked you.

OpenStudy (mathmale):

Following Solomon's advice: Identify all the possible integer factors of 24; for example: 1, 2, 3, 4, 6, 8, .... Next, if you know synthetic division, by all means check these factors one by one to determine which among them leave no remainder; any such factor of 24 is a root of this polynomial. If you don't know synth. div., well, you have more work to do. If you want to check whether 2 is a root (for example), convert 2 to a binomail possible factor: (x-2), and divide that into the given polynomial.

OpenStudy (solomonzelman):

mathmale, I don't think that this question is from college algebra, although it is a cubic function. (nothing as bad as f(x)=x^3+x+31 :) )

OpenStudy (solomonzelman):

Well, either way... but I wouldn't take a harder approach, when I can take an easier one.

OpenStudy (mathmale):

Hi, Solo...which approach do you think is the harder one?

OpenStudy (solomonzelman):

Factoring is right there... whereas, all possible zeros; plugging them in and seeing which work? Well, you would finish in a second doing either of the approaches, but I am I probably the laziest user on here, so I prefer...

hero (hero):

I prefer factoring by grouping for this one: x^3 - 2x^2 + 16x - 32 = x^2(x - 2) - 16(x - 2)

OpenStudy (solomonzelman):

._.

OpenStudy (mathmale):

I agree that factoring by grouping would likely be the fastest approach here. @millalie94: would you please continue the factoring by grouping begun by Hero? Ask questions about this if you need to.